If 50 is decreased by 4times a certain number, the result is 15 more than the number. Find the number
Answer by macston(5194)

.
X=the number
.
50-4X=X+15
35=5X
7=X
ANSWER: The number is 7.
.
CHECK:
50-4X=X+15
50-4(7)=7+15
50-28=22
22=22